9. Find AG at 37 C for a reaction given the following: 4,Sº= -52.9 J/Kmol and AHº= -52.9 kJ/mol
Answer #1

oce = BH - TDS. we have given DH = -52.9 KT/mol. As = 52.9 / le mol In 37°C 2 37 +273 = 310k. we can find see by es su = (-52